4. The threshold frequency of a metal is 4 × 10 14 s –1 . The minimum energy which photons must possess to produce photoelectric effect with the metal is (h = 6.6 × 10 –34 Js)

4.     The threshold frequency of a metal is 4 × 1014 s–1. The minimum energy which photons must possess to produce photoelectric effect with the metal is (h = 6.6 × 10–34 Js)

Grade:11

1 Answers

Khandavalli Satya Srikanth
10 years ago

Minimum energy required to produce photoelectric effect is threshold energy which is given by

E=h*frequency

E=(6.6*10^-34)*(4*10^14)=2.64*10-19 J